OpenStudy (anonymous):

Can someone explain to me how I can differentiate this equation?: f(x) = x^2 * e^-x

OpenStudy (anonymous):

use product rule

OpenStudy (anonymous):

product went work try quetient rule

OpenStudy (anonymous):

ur answer is (2x-x^2)/e^x

OpenStudy (anonymous):

use \[(fg)'=f'g+g'f\] with \[f(x)=x^2, f'(x)=2x,g(x)=e^{-x},g'(x)=-e^{-x}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Just, why did the \[e ^{-x}\] get: \[-e^{-x}\] I'm kinda new to calculus.